OpenStudy (loser66):

Prove \(3^{2^n}-1\) divisible by \(2^{n+2}\). Please, help

OpenStudy (loser66):

I tried and it is a mess!!! :)

OpenStudy (loser66):

Likely we need prove \(3^{2^{n+1}}-1\) divided by \(2^{n+1}\) given that \(3^{2^n}-1\) divided by \(2^n\) it is an ugly step.

OpenStudy (loser66):

I tried: \(2^n\) is an even number, hence, it can be written as 2m. Then \(3^{2m}-1= 9^m-1=(9-1)(9^{m-1}+9^{m-2}+\cdots +1)\) \(9-1) =8 \div 4\), hence the original one divided by \(4=2^2\) The leftover is to prove it is divided by \(2^n\) but how??

OpenStudy (holsteremission):

Assume the statement holds for \(n=k\), then use that fact to show it holds for \(n=k+1\). \[3^{2^{k+1}}-1=\left(3^{2^k}\right)^2-1=\left(3^{2^k}-1\right)\left(3^{2^k}+1\right)\]By the hypothesis, you know that \(3^{2^k}-1\) is divisible by \(2^{k+2}\), so all you need to do is show that \(3^{2^k}+1\) is divisible by \(2\). This could probably also be shown inductively, but I don't see a need for that unless you want to be as explicit as possible in showing that every power of \(3\) is odd.

OpenStudy (loser66):

Thanks a ton @HolsterEmission I got it. It is elegant logic when you turn \(3^{2^{k+1}}\) to \((3^{2^k})^2\)